Sara spends her entire weekly food allowance of $42 on Zinger and soft drinks. The price of a Zinger is $2, and the price of a soft drink is $1. Sara purchases 12 Zingers and 18 soft drinks, and her marginal rate of substitution between Zingers and soft drinks is 1. Is this the optimal consumption of Sara? Explain.
1
Expert's answer
2014-11-17T12:37:11-0500
B = $42, Pz = $2, Psd = $1. Qz = 12 Zingers, Qsd = 18 soft drinks, MRSz sd = 1.
As MRSxy = MUx/MUy, then MUz/MUsd = 1, so MUz = MUsd
The optimal consumption that maximizes utility is in the point, where MUx/Px = MUy/Py, but in our case as MUz = MUsd, so MUz/$2 is not equal to MUsd/$1.
That's why this is not the optimal consumption of Sara and she needs to decrease consumption of Zinger and increase consumption of soft drinks, until MUz/Pz = MUsd/Psd.